A metal aphere of radius R is charged to a potential V. Find the electrostatic energy stored in the electric field within a concentric sphere of radius 2R.

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

V=q/(4piε0R), strength of electric field outside metal sphere q/(4piε0r),r is distance from centre. energy density of electric field=

(1/2)ε0E2, total energy=∫(1/2)ε0(VR/r)2 dr from r=R to r=2R